A consumer affairs investigator records the repair cost for 1010 randomly selected VCRs. A sample mean of $72.82 and standard deviation of $24.17 are subsequently computed. Determine the 90% confidence interval for the mean repair cost for the VCRs. Assume the population is approximately normal.

1: Find the critical value that should be used in constructing the confidence interval. Round your answer to three decimal places.

2. Construct the 90% confidence interval. Round your answer to two decimal places.
Answer #1

Answer

(A) given that we have to find 90% confidence interval

so. alpha = 1- confidence level

= 1-0.90

= 0.10

and

degree of freedom = n-1

= 10-1

= 9

t critical = T.INV.2T(alpha,df)

= T.INV.2T(0.10,9)

= 1.833

(B) using TI 84

press stat then tests then TInterval

enter data

x = 72.82

s = 24.17

n = 10

c-level = 0.90

press enter, we get